A point charge of 16 uc is placed in the centre of a conductive shell carrying a charge of -8 uc. What are the charges on the inner and outer surfaces of the shell?


1
Expert's answer
2022-03-01T15:23:16-0500

Now surface charge density

"\sigma_a=\frac{Q}{A}\\\sigma_a=\frac{-8\times10^{-6}}{4\pi a^2}"

Now

"\sigma_b=\frac{q+Q}{A}"

"\sigma_b=\frac{(16-8)\times10^{-6}}{4\pi b^2}"

"\sigma_b=\frac{(8)\times10^{-6}}{4\pi b^2}"